If sin [cot–1 (x + 1)] = cos(tan–1x), then find x.this question solution please.

Grade:12th pass

1 Answers

User
7 years ago
sin(cot^-1(x+1)=cos(tan^-1(x)-->sin(sin^-1(1/sqrt(x^2+2x+2)=cos(cos^-1(1/sqrt(1+x^2)--> 1/sqrt(x^2+2x+2)=1/sqrt(1+x^2)Solving the above equation we get:x=-1/2